# Gradient Extrapolation for Debiased Representation Learning

### v1

@@ Appendix F Implementation Details
F.1 Implementation details on Datasets-1
@@
For the C-MNIST [3, 27], we deploy a multi-layer perceptron (MLP) with three fully connected layers, while for C-CIFAR-10 [15, 31] and bFFHQ [22, 27], we employ a pretrained ResNet-18 model [14] (pretrained on ImageNet1K [8]) as the backbone. The Stochastic Gradient Descent (SGD) optimizer, with a momentum of 0.9 and a weight decay of 0.01, is applied across all three datasets. Batch sizes are configured as follows: 100 per group/pseudo-group for C-MNIST and C-CIFAR-10, and 32 for bFFHQ. Learning rates are set to 0.1 for C-MNIST in the known attribute case and 0.01 in the unknown attribute case. For C-CIFAR-10 and bFFHQ, the learning rate is set to 0.0001.
@@
For GERNE in the known attribute case, we present results from two experimental configurations. In the first experiment, we set $c=1\text{ and }\beta=0$, which corresponds to resampling [16] from groups, i.e., training on $\mathcal{L}_{lb}$ without extrapolation. In the second experiment, $\beta$ is tuned for $c\in\{\frac{1}{2},1\}$. In the unknown attribute case, $t$ is an additional hyperparameter to be tuned. We avoid using any data augmentations as certain transformations can unintentionally fail to preserve the original label. For example, flips and rotations in C-MNIST can distort labels (e.g., a rotated ”6” appearing as a ”9”) [41]. For training $\tilde{f}$ in case of unknown attributes in the training set, we employ the same model architecture as $f$, with modifications to the hyperparameters: the weight decay is doubled, and the learning rate is reduced to one-tenth of the learning rate used to train $f$. The loss function used is the cross-entropy loss in all the experiments.
F.2 Implementation details on Datasets-2
@@
To ensure a fair comparison of GERNE with other methods in [49], we adhere to the same experimental settings. For the Waterbirds [45] and CelebA [30] datasets, we utilize a pretrained ResNet-50 model [14] as the backbone, while for CivilComments [5], we use a pretrained BERT model [9]. Each backbone is followed by an MLP layer with $K$ output neurons. We employ SGD with a momentum of 0.9 and a weight decay of 0.01 for Waterbirds and CelebA, while for CivilComments, we use AdamW [24] optimizer with a weight decay of 0.0001 and a tunable dropout rate. We set batch sizes to 32 for both Waterbirds and CelebA and 5(16) per group(pseudo-group) for CivilComments. The learning rates are configured as follows: 0.0001 for Waterbirds and CelebA, and 0.00001 for CivilComments. Additionally, we set the bias reduction factors $c$ to 0.5 for Waterbirds and CelebA and to 1 for CivilComments. For image datasets, we resize and center-crop the images to 224×224 pixels. In the case of unknown attributes in the training set, $\tilde{f}$ has the same architecture as $f$, but we adjust the hyperparameters: the weight decay is doubled, and the learning rate is reduced to one-tenth of the value used to train $f$. We employ the Cross-entropy loss as the loss function across all experiments. For experiments with unknown attributes in both the training and validation sets, we limit the hyperparameter $t$ search space to the interval $[0,\frac{1}{2}]$
Appendix G Evaluating GERNE under limited attribute information
@@
To further demonstrate the effectiveness of GERNE in scenarios with limited access to samples with attribute information, we conduct two experiments on the CelebA dataset. In these experiments, we exclude the training set and only use the validation set with its attribute information for training. We follow the same settings and implementation details described earlier. As part of the implementation, we first tune the hyperparameters using the designated evaluation metric. Once we determine the optimal hyperparameters, we fix them and train the model $f$ three times with different random seeds. Finally, we report the average worst-group test accuracy and standard deviation across these runs.
@@
Experiment 1 - Evaluation on test set. In this experiment, $f$ is trained using the full validation set. The worst-group accuracy on the test set is used as the evaluation metric. This setup represents the best possible performance achievable when relying solely on the validation set for training.
@@
Experiment 2 - Cross-validation. we divide the validation set into three non-overlapping folds, ensuring that each fold preserves the same group distribution as the original validation set (i.e., we randomly and equally distribute samples from each group across the folds). We use two folds to train $f$ and reserve the remaining fold for hyperparameter tuning and model selection, using the worst-group accuracy on this fold as the evaluation metric. We repeat this process three times, with each fold serving as the validation fold exactly once. We summarize the average worst-group test accuracy and standard deviation across all nine runs (three folds × three seeds) in Tab. 3.
@@
We also compare these results with DFR, a method that trains the last layer of the model on the validation set after performing ERM training on the training set. GERNE consistently achieves state-of-the-art results, demonstrating its robustness and effectiveness even with severely limited attribute information.
Table 3: Performance Comparison of GERNE and DFR using the validation set for training.
@@
GERNE - Evaluation on test set 90.97 $\pm$ 0.35
GERNE - Cross-validation 88.63 $\pm$ 0.59

### v2

@@ Appendix F Implementation Details
F.1 Implementation Details for Datasets-1
@@
For the C-MNIST [3, 27], we deploy a multi-layer perceptron (MLP) with three fully connected layers, while for C-CIFAR-10 [15, 31] and bFFHQ [22, 27], we employ ResNet-18 model [14], pretrained on ImageNet1K [8], as the backbone. We apply the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 and a weight decay of $10^{-2}$ across all three datasets. We set the batch size to 100 per group or pseudo-group for both C-MNIST and C-CIFAR-10, and to 32 for bFFHQ. For C-MNIST, we use a learning rate of $10^{-1}$ in the known attributes case and $10^{-2}$ in the unknown attributes case. For C-CIFAR-10 and bFFHQ, we use a learning rate of $10^{-4}$.
@@
In the unknown attribute case, we treat the threshold $t$ as an additional hyperparameter that requires tuning. We avoid using any data augmentations, as certain transformations can unintentionally fail to preserve the original label. For example, flips and rotations in C-MNIST can distort labels (e.g., a rotated “6” appearing as a “9”) [41]. For training $\tilde{f}$ in case of unknown attributes in the training set, we employ the same model architecture as $f$, with modifications to the hyperparameters: the weight decay is doubled, and the learning rate is reduced to one-tenth of the learning rate used to train $f$. The loss function used is the Cross-entropy loss for all the experiments.
F.2 Implementation Details for Datasets-2
@@
To ensure a fair comparison between GERNE and the methods in [50], we adopt the same experimental settings. For Waterbirds [46] and CelebA [30] datasets, we use a pretrained ResNet-50 model [14] as the backbone, while for CivilComments [5], we use a pretrained BERT model [9]. We append an MLP classification head with $K$ outputs. We employ SGD with a momentum of 0.9 and a weight decay of $10^{-2}$ for Waterbirds and CelebA. For CivilComments, we use AdamW [24] optimizer with a weight decay of $10^{-4}$ and a tunable dropout rate. We set batch sizes to 32 for both Waterbirds and CelebA and 5(16) per group(pseudo-group) for CivilComments. The learning rates are configured as follows: $10^{-4}$ for Waterbirds and CelebA, and $10^{-5}$ for CivilComments. Additionally, we set the bias reduction factors $c$ to 0.5 for Waterbirds and CelebA and to 1 for CivilComments. For image datasets, we resize and center-crop the images to 224×224 pixels. In the case of unknown attributes in the training set, $\tilde{f}$ has the same architecture as $f$, but we adjust the hyperparameters: the weight decay is doubled, and the learning rate is reduced to one-tenth of the value used to train $f$. We employ the Cross-entropy loss as the loss function across all experiments. For experiments with unknown attributes in both the training and validation sets, we limit the search space for $t$ to the interval $0,\frac{1}{2}]$.
Appendix G Evaluating GERNE with Limited Attribute Information
@@
To further demonstrate the effectiveness of GERNE in scenarios with limited access to samples with known attributes, we conduct two experiments on the CelebA dataset using only the validation set with its attribute information for training (excluding the training set). We follow the same experimental setup and implementation details outlined in Sec. F.2. As part of the implementation, we first tune the hyperparameters using the designated evaluation metric. Once we determine the optimal hyperparameters, we fix them and train the model $f$ three times with different random seeds. Finally, we report the average worst-group test accuracy and standard deviation across these runs.
@@
Evaluation on the Test Set. In this experiment, we train the model $f$ using the entire validation set and use the worst-group test accuracy as the evaluation metric. This setup represents the best possible performance achievable when relying solely on the validation set for training.
@@
Cross-Validation. In this experiment, we divide the validation set into three non-overlapping folds, ensuring that each fold preserves the same group distribution as the original validation set. Specifically, we randomly and evenly distribute samples from each group in the validation set across the folds. We train $f$ using two of the folds, and use the remaining fold for hyperparameter tuning and model selection, where the worst-group accuracy on this fold serves as the evaluation metric. We repeat this process three times so that each fold is used once as the validation fold. Finally, we report the average worst-group test accuracy and standard deviation across all nine runs (three folds × three seeds) in Tab. 3.
@@
We compare the results of GERNE with DFR, a method that trains the final layer on the validation set following ERM training on the training set. GERNE consistently achieves state-of-the-art results, demonstrating its robustness and effectiveness even under limited attribute information.
Table 3: Performance comparison of GERNE and DFR. GERNE uses only the validation set for training. We report the worst-group test accuracy (%) and standard deviation over three trials.
@@
GERNE — Evaluation on Test Set 90.97 $\pm$ 0.35
GERNE — Cross-Validation 88.63 $\pm$ 0.59